OpenStudy (anonymous):

Which is the exact value of the expression √48 - √75 + √192?

OpenStudy (anonymous):

12.23218924

OpenStudy (mertsj):

\[\sqrt{48}-\sqrt{75}+\sqrt{192}=\sqrt{16(3)}-\sqrt{75(3)}+\sqrt{64(3)}=\]

OpenStudy (mertsj):

\[\sqrt{16}\sqrt{3}-\sqrt{25}\sqrt{3}+\sqrt{64}\sqrt{3}=4\sqrt{3}-5\sqrt{3}+8\sqrt{3}=7\sqrt{3}\]

OpenStudy (anonymous):

Thank you!